### v1

@@ 4.1 RQ1: How do open-source projects structure and document their PTM dependencies?
To answer RQ1, we analyze both the structure of dependencies and their documentation. We design a multi-step analysis with shared procedures to address both aspects. This analysis leverages 401 projects processed as described in §3.2.
@@ 4.1.1 Methodology
Step 1: PTM Model Name Identification.
For both sub-RQs, we need to identify the PTMs used in target projects. We begin with the model-loading signatures identified by the PeaTMOSS dataset. PeaTMOSS provides direct mappings between known loading functions and PTM names when those names are statically defined, i.e., hardcoded string literals in the source code. These mappings are already included in the dataset. For example:
  • torchvision.models.resnet18(pretrained=True) (TorchVision)
In contrast, for dynamically loaded models, names are not fixed in the code but stored in variables, passed as arguments, or read from configuration files or command-line inputs. Although PeaTMOSS includes the loading function (e.g., from_pretrained(...)), it does not identify the exact model name. We start from the files where these calls appear and manually check nearby code and other project files collected in Section 3.2 to manually trace the data flow and find the actual model names being loaded. In total, we identify 1059 unique PTMs loaded across the 401 projects. Figure 5(a) shows the total number of PTMs by model type across all projects. NLP models are the most prevalent (169 PTMs), followed by vision models (136 PTMs). Multi-modal models are also substantial, with 51 PTMs, while speech- and audio-focused models are less common, with 20 and 5 PTMs, respectively.
Step 2: Manual Examination of PTM Dependency Structure (for RQ1.1).
@@
Using the PTM names extracted in Step 1, we first classify each repository as using either a single PTM or multiple PTMs. For repositories containing multiple PTMs, we examine the relationships between them at the project level, specifically asking whether the PTMs used within a project belong to the same base family or span across different families. To support this analysis, we group individual PTMs into broader base model families, where each family represents a core architecture shared by several model variants. A model family groups PTMs with the same underlying architecture. For example, the BERT family includes bert-base-uncased, roberta-base, and distilbert-base-uncased. This approach aligns with prior literature treating models such as RoBERTa and DistilBERT as BERT variants (Liu et al., 2019; Sanh et al., 2019).
@@
We assign models to (base model) families using a two-step process. First, we apply a keyword-based heuristic that relies on architecture-identifying substrings in model names. For instance, any model whose name contains “bert” was assigned to the BERT family, “gpt” to GPT. This approach is supported by prior work (Jiang et al., 2023a; Hao et al., 2023), which reports that Hugging Face model names generally reflect underlying architectures. Second, when model names do not explicitly include the family name, we manually map them to the appropriate family. For example, inpainting models such as anything-4.0-inpainting, which omit “Stable Diffusion” but are known community variants, Stable Diffusion family based on official documentation (Hugging Face, 2021; PyTorch, 2023) and our prior knowledge. Labeling is performed by one author and fully reviewed by another. Inter-annotator agreement was 85%, and disagreements are resolved through discussion. Because architecture is typically embedded in model names, this task involves little subjectivity.  [^3] Note that Hugging Face provides metadata fields such as model_type for each model hosted on the hub. However, this metadata does not consistently capture shared architectures and is therefore unreliable for family-level grouping (Jiang et al., 2025).
@@
Step 3: Identify PTM Name Declarations (for RQ1.2).
@@
Step 4: Identify PTM Version Declarations (for RQ1.2).
Versioning promotes the reproducibility, stability, and traceability of machine learning experiments. Since pretrained models may change over time (e.g., due to updated weights, bug fixes, or architectural changes), specifying the version of a PTM used in a project helps guarantee consistent behavior across environments and over time.
@@
To identify version declarations, we statically analyze model-loading code and check whether these version parameters (e.g., revision in Hugging Face, version tags in PyTorch Hub) are explicitly provided. This allows us to measure how often projects declare exact PTM versions, and thus, how they approach reproducibility in their dependency practices.
@@ 4.1.2 Results
PTM Dependency Structure. Our findings reveal two high-level PTM usage patterns: Single PTM and Multiple PTMs. Figure 5(b) summarizes the distribution of target projects, grouped by their model type. Most single-PTM projects rely exclusively on vision models (113 projects) or NLP (55 projects), with only a handful centered on speech, audio, or multi-modal models. By contrast, multi-PTM projects overwhelmingly involve cross-modality combinations, most commonly NLP + vision (8 projects). Other combinations, such as multi-modal + vision (8 projects), multi-modal + NLP (3 projects), or NLP + speech (3 projects), are less frequent but illustrate the diversity of integration patterns. Multiple-PTM projects tend to combine models across different modalities.
(a) Counts of individual PTMs by model type.
(b) Counts of projects by PTM type combinations.
Figure 5: Distribution of PTM usage: (a) number of models by type and (b) number of projects by modality combinations.
@@
PTM Usage in OSS ProjectsSingle PTM(48%)Multiple PTMs(52%)Interchangeable(37%)Complementary(23%)
Figure 6: PTM Dependency Structures in OSS Projects. Some projects include both interchangeable and complementary PTMs, which is why the percentages add up to more than 52%.
@@
In our dataset, interchangeable PTMs are typically drawn from the same family, while complementary PTMs typically combine models from different families to address different tasks or modalities. This observation is based on a manual review of 50 multi-PTM projects, of which 81% used PTMs from the same family in interchangeable roles and 85% combined PTMs from different families in complementary roles. Exceptions exist: some projects employ multiple models from the same family for different tasks, while others use models from different families to perform overlapping or similar functions. Thus, although this heuristic reflects common patterns in PTM usage, it should not be interpreted as a universal rule.
These two dependency structures reflect different design goals: interchangeable PTMs offer flexibility and configurability, while complementary PTMs enable functional decomposition and complex, modular architectures. These findings highlight how developers balance flexibility and specialization: interchangeable PTMs allow runtime configurability within a single family, while complementary PTMs enable modular, multi-modal, or cross-stage workflows across families. A comparison of their characteristics is presented in Table 2.
@@
PTM Dependency Documentation. Our analysis revealed three high-level PTM dependency patterns. The first pattern is projects with no PTM declarations outside of the source code. The second pattern includes projects with partial or scattered mentions. The third pattern covers projects with explicit and centralized documentation.
Among the projects, 58.9% (236 projects) mention PTMs only in source code and not in configuration or documentation files. Identifying the actual PTM dependencies in these cases requires code analysis, especially for models loaded dynamically at runtime. Another 20.0% (80 projects) provide partial mentions of PTMs. Figure 7 details this breakdown, showing that 12.0% (48 projects) mention PTMs only in documentation, 3.5% (14 projects) only in configuration files, and 4.5% (18 projects) in both. These mentions are often scattered or informal, such as listing model names without version information or source specification. Only 21.2% (85 projects) explicitly declare all used PTMs in external files. Of these, 15.5% (62 projects) do so only in documentation, 1.5% (6 projects) only in configuration files, and 4.2% (17 projects) in both. Our review also found that even when PTM names are present in configuration files (e.g., YAML or JSON), they rarely include explicit hub information, such as Hugging Face or PyTorch Hub, which limits reproducibility and automation.
@@
4.2 RQ2: What are the stages and their organization in the pipeline of PTM-based open source projects?
@@ 4.2.1 Methodology
The search is conducted across the following databases: IEEE Xplore, ACM Digital Library, and Google Scholar (queried using the Publish or Perish tool. [^4] See https://harzing.com/resources/publish-or-perish.) After removing duplicates and irrelevant papers based on titles and abstracts, 300 scholarly papers are selected for full-text review.
Our review finds no prior work provides a taxonomy of PTM reuse pipelines or their stages. Most papers describe isolated PTM reuse activities (e.g., fine-tuning, prompt engineering) without situating them in a unified pipeline structure. Based on our analysis, we manually identify a minimal, reusable set of PTM-centric stages that consistently appear across both the literature and open-source projects. We focus specifically on model-level, code-observable practices, and intentionally exclude full-stack, DevOps, or governance frameworks (Fitsilis et al., 2024). Table 3 lists the eight identified stages. Stages such as data processing and post-processing are treated at a coarse level, as their fine-grained substeps (e.g., data acquisition, data preparation) vary widely and are often implicit or domain-specific.
@@
While the first author is primarily responsible for labeling the pipeline stages across all target projects, the two authors participate in multiple rounds of collaborative discussion to refine the stage definitions. To validate the reliability of the coding scheme, we conduct a cross-check by having a third author independently annotate a randomly selected subset of 282 PTMs. This subset is generated using a random sampling approach calculated to achieve a 95% confidence level and a 5% margin of error, ensuring it is statistically representative of the full dataset. The comparison between the two annotators yields an inter-rater agreement score of 0.81, indicating substantial agreement (Viera et al., 2005).
@@ 4.2.2 Results
  1. 1.
    Evaluation (EVAL) assesses model outputs systematically using quantitative metrics (e.g., FID score, BLEU, accuracy) or qualitative criteria to measure quality and task alignment. In practice, evaluation is often coupled with validation testing and feedback loops, which provide continuous performance checks, guide hyperparameter adjustments, and prevent overfitting during model adaptation.
  2. 2.
Table 3: Unified taxonomy of PTM reuse pipeline stages from our empirical study of 401 projects, arranged chronologically from model initialization to final delivery. Stages that differ from or expand on prior literature are highlighted in bold, including Adaptation (ADPT), Prompt Generation (PROM), Post-Processing (POST), and the newly surfaced Evaluation (EVAL) and Delivery (DLV) stages. Literature-based highlights and our taxonomy-based descriptions are provided for each stage.
@@
Prompt Generation (PROM) Crafting prompts—manually or automatically—to guide generative PTMs (Ramachandran, 2023). Creates prompts manually, templated, or automatically using prompt engineering techniques.
@@
Using PTM Weights Only the PTM’s weights are used to initialize a different model structure.
Example: Loading BERT weights into a variant with an added adapter layer.
Addition A new prediction head is appended to the base PTM, which remains otherwise unchanged.
@@
In addition, several existing stages take on different emphases in practice. For example, Adaptation appears prominently in real-world pipelines, encompassing not only the addition of a head but also head replacement, selective layer use, architectural changes, and configuration tweaks (Table 4). Fine-tuning itself spans a spectrum of strategies, from fully updating all parameters to partially freezing layers (e.g., freezing lower-level feature extractors while training the higher-level layers), thereby balancing computational efficiency with task specialization. Likewise, Prompt Generation and Post-Processing often serve as critical operational stages in generative tasks, with the latter ensuring output quality and safety.
@@
Feature ExtractionGenerativeDiscriminativeModel InitializationModel InitializationModel InitializationModel AdaptationModel AdaptationModel AdaptationData ProcessingData ProcessingData ProcessingPrompt GenerationFeature EngineeringFeature EngineeringFeature EngineeringFine-TuningFine-TuningFine-TuningInferenceInferenceInferencePost-ProcessingPost-ProcessingPost-ProcessingEvaluationEvaluationDeliveryDelivery
Figure 8: Visualization of three pipeline organizations for reusing pretrained models. Grey boxes indicate stages common to most pipelines, i.e., Fine-Tuning and Inference. Dashes indicate optional stages.
@@
    Generative-Oriented Pipeline. This organization centers on generating novel outputs such as text, code, or images. After model initialization, adaptation, and data processing, a defining stage is prompt generation, which guides the model’s generative behavior. Depending on the application, the PTM may be used as-is via inference, or fine-tuned for greater task alignment. Feature engineering may be incorporated in hybrid methods (e.g., retrieval-augmented generation). As in other organizations, optional stages such as post-processing and delivery serve to clean, format, or serve the outputs.
@@
RQ2: What are the stages and their organization in the pipeline of PTM-based open source projects? Our analysis reveals that PTM reuse pipelines follow three dominant organizations: feature extraction, generative, and discriminative. Across all types, inference is the most consistently used stage, underscoring its central role in practical deployments. While some applications adopt minimal adaptation, complex workflows incorporate full-stage pipelines, including data processing, fine-tuning, post-processing, and delivery. The Model Adaptation stage is especially diverse; developers reuse PTMs as-is, initialize with pre-trained weights, add task-specific heads, or selectively modify layers. These patterns demonstrate that PTM reuse is rarely plug-and-play; instead, customization is key to optimizing performance and aligning with task-specific requirements across varied real-world scenarios.

### v2

@@ 4.1 RQ1: How do open-source projects structure and document their PTM dependencies?
To answer RQ1, we manually analyze both the structure of dependencies and their documentation. This analysis leverages 401 projects processed as described in §3.2.
@@ 4.1.1 Methodology
Step 1: Manual Examination of PTM Dependency Structure (for RQ1.1).
@@
Using the PTM names extracted in 3.2, we first classify each repository as using either a single PTM or multiple PTMs. For repositories containing multiple PTMs, we examine the relationships between them at the project level, specifically asking whether the PTMs used within a project belong to the same base family or span across different families. To support this analysis, we group individual PTMs into broader base model families. We define a base model family as a group of pretrained models that share the same underlying neural architecture and primary pretraining objective, while differing in training data, scale, or optimization strategy. This definition captures models that are functionally similar in their learned representations despite variations in training methodology. For instance, the BERT family includes bert-base-uncased, roberta-base, distilbert-base-uncased, and albert-base-v2. Compared to the original BERT model, RoBERTa modifies the optimization and training configuration, DistilBERT reduces model size via distillation, and ALBERT introduces parameter sharing. Despite these differences, all retain the bidirectional transformer encoder with masked language modeling pre-training and are therefore considered part of the same base model family. This categorization aligns with the existing literature that treats models such as RoBERTa and DistilBERT as BERT variants (Liu et al., 2019; Sanh et al., 2019).
@@
We assign models to (base model) families using a two-step process. First, we apply a keyword-based heuristic that relies on architecture-identifying substrings in model names. For instance, any model whose name contains “bert” was assigned to the BERT family, “gpt” to GPT. This approach is supported by prior work (Jiang et al., 2023a; Hao et al., 2023), which reports that Hugging Face model names generally reflect underlying architectures. Second, when model names do not explicitly include the family name, we manually map them to the appropriate family. For example, inpainting models such as anything-4.0-inpainting, which omit “Stable Diffusion” but are known community variants, Stable Diffusion family based on official documentation (Hugging Face, 2021; PyTorch, 2023), and our prior knowledge. One author performed the initial labeling, and a second author independently reviewed all labels as a validation step. The percent agreement between the reviewer and the original annotations was 85%. All discrepancies were discussed and resolved. Note that Hugging Face provides metadata fields such as model_type for each model hosted on the hub. However, this metadata does not consistently capture shared architectures and is therefore unreliable for identifying model families (Jiang et al., 2025).
@@
Step 2: Identify PTM Name Declarations (for RQ1.2).
@@
Step 3: Identify PTM Version Declarations (for RQ1.2).
Versioning promotes the reproducibility, stability, and traceability of ML experiments. Since pretrained models may change over time (e.g., due to updated weights, bug fixes, or architectural changes), specifying the version of a PTM used in a project helps guarantee consistent behavior across environments and over time.
@@
To identify version declarations, we statically analyze model-loading code and check whether these version parameters (e.g., revision in Hugging Face, version tags in PyTorch Hub) are explicitly provided. In addition to source code, we also analyze configuration files (e.g., YAML and JSON files) and project README files to identify explicit references to PTM versions. This allows us to measure how often projects declare exact PTM versions, and thus, how they approach reproducibility in their dependency practices.
@@ 4.1.2 Results
Prevalence of Model Families. The sampled 401 projects exhibit several well-represented model families rather than a single dominant group. Across the 381 unique PTMs identified, we find 79 base model families. The most represented family is BERT (12.1%, 46/381), including models such as bert-base-uncased, roberta-base, and distilbert-base-uncased. ResNet follows with 10.5%(40/381), including resnet50, resnet101, and resnext50_32x4d. ViT accounts for 9.2% (35/381), with examples such as vit_base_patch16_224, ViT-B-16, and dino_vitb16. Stable Diffusion contributes 7.1% (27/381), including stable-diffusion-v1-5 and stable-diffusion-xl-base-1.0, while the GPT family represents 5.5% (21/381), including gpt2, gpt2-large, and EleutherAI/gpt-neo-2.7B. Other notable families include CLIP (4.5%), T5 (4.2%), Wav2Vec 2.0(3.4%), VGG (2.4%), BART (2.1%), and MobileNet (1.6%). Overall, the distribution reflects a broad and diverse ecosystem spanning vision, language, speech, and generative models.
PTM Dependency Structure. To characterize how OSS projects organize their PTM dependencies, we consider the number of PTMs integrated within a project (single vs. multiple) and the modality and architectural type of the models (base model family) involved. These dimensions jointly capture structural complexity at different levels. Figure 6 summarizes the distribution of target projects, grouped by their model type. Most single-PTM projects rely exclusively on vision models (113 projects) or NLP (55 projects), with only a handful centered on speech, audio, or multi-modal models. By contrast, multi-PTM projects overwhelmingly involve cross-modality combinations, most commonly NLP + vision (8 projects). Other combinations, such as multi-modal + vision (8 projects), multi-modal + NLP (3 projects), or NLP + speech (3 projects), are less frequent but illustrate the diversity of integration patterns. Multiple-PTM projects tend to combine models across different modalities.
Figure 6: Number of projects by modality combinations.
@@
PTM Usage401 projectsSingle PTM193 (48%)Multiple PTMs208 (52%)Interchangeable149(37% of all 401)Complementary92(23% of all 401)
Figure 7: PTM dependency structures in OSS projects. The top-level percentages (48%, 52%) represent proportions across all 401 projects. For each subcategory, we report both the count and its percentage relative to the full set of 401 projects (rather than only multi-PTM projects). Some multi-PTM projects contain both interchangeable and complementary PTMs; therefore, the subcategory totals are not mutually exclusive (149 + 92 >> 208).
@@
In our dataset, interchangeable PTMs are typically drawn from the same family, while complementary PTMs typically combine models from different families to address different tasks or modalities. This observation is supported by a manual review of 135 randomly sampled multi-PTM projects out of 208 multi-PTM projects, where the sample size was determined to achieve a 95% confidence level with a 5% margin of error. Our results show that 76% of the projects used PTMs from the same family in interchangeable roles and 86% of the projects combined PTMs from different families in complementary roles. Despite these dominant patterns, the mapping between family membership and functional role is not strict. Some projects use models from different families interchangeably for the same task, while others employ multiple models from the same family in complementary roles. For instance, UMR [^3] https://github.com/NVlabs/UMR implements a configurable perceptual network that accepts vgg16, alexnet, resnet, or squeezenet as a runtime parameter: all serving the identical role of computing perceptual similarity, despite belonging to different families. Conversely, lama-cleaner [^4] https://github.com/sanster/lama-cleaner uses multiple Stable Diffusion variants not as interchangeable alternatives, but in distinct stages of its image inpainting pipeline.
These two dependency structures reflect different design goals: interchangeable PTMs typically offer runtime flexibility and configurability, often within the same family but occasionally across families; complementary PTMs typically enable functional decomposition through modular, multi-modal, or multi-stage architectures, often across families but occasionally within one. A comparison of their typical characteristics is presented in Table 2.
@@
PTM Dependency Documentation. Our analysis revealed three high-level PTM dependency patterns. The first pattern is defined by projects with no PTM declarations outside of the source code. The second pattern includes projects with partial or scattered mentions. The third pattern covers projects with explicit and centralized documentation.
Among the projects, 58.9% (236 projects) mention PTMs only in source code and not in configuration or documentation files. Identifying the actual PTM dependencies in these cases requires code analysis, especially for models loaded dynamically at runtime. Another 20.0% (80 projects) provide partial mentions of PTMs. Figure 8 details this breakdown, showing that 12.0% (48 projects) mention PTMs only in documentation, 3.5% (14 projects) only in configuration files, and 4.5% (18 projects) in both. These mentions are often scattered or informal, such as listing model names without version information or source specification. Only 21.2% (85 projects) explicitly declare all used PTMs in external files. Of these, 15.5% (62 projects) do so only in documentation, 1.5% (6 projects) only in configuration files, and 4.2% (17 projects) in both.
Our review also found that even when PTM names are present in configuration files (e.g., YAML or JSON), they rarely include explicit hub information, such as Hugging Face or PyTorch Hub, which limits reproducibility and automation.
@@
Documentation Practices in Single-PTM vs. Multi-PTM Projects. To understand how project complexity affects documentation practices, we analyzed whether single-PTM and multi-PTM projects differ in their dependency documentation approaches. Among single-PTM projects, we found that 70.5% (136/193) of single-PTM projects mention their PTM only in source code, while 29.5% (57/193) explicitly declare it in outside source code files: 20.7% (40 projects) in README only, 2.6% (5 projects) in configuration files only, and 6.2% (12 projects) in both.
In contrast, multi-PTM projects introduce a third category: partial documentation, where only some PTM dependencies are documented outside the source code. Among multi-PTM projects, 48.1% (100/208) mention PTMs only in source code, and 38.5% (80/208) provide partial mentions: 23.1% (48 projects) in README only, 6.7% (14 projects) in configuration files only, and 8.7% (18 projects) in both. Only 13.5% (28/208) explicitly declare all their PTM dependencies in external files: 8.2% (17 projects) in README only, 0.5% (1 project) in configuration files only, and 4.8% (10 projects) in both. While multi-PTM projects more frequently begin documenting their dependencies outside source code (51.9% vs. 29.5%), they rarely achieve complete coverage (13.5% vs. 29.5%), suggesting that increasing project complexity motivates documentation efforts but simultaneously makes comprehensive documentation harder to maintain.
@@
4.2 RQ2: What are the stages and their organization in the reuse pipeline of PTM-based open source projects?
@@ 4.2.1 Methodology
We derived the search keywords from established ML workflow terminology (e.g., pipeline, workflow, lifecycle) together with PTM-specific concepts (e.g., pre-trained model, foundation model, large language model). During the scoping phase, we evaluated additional candidate terms, such as “transfer learning pipeline” and “fine-tuning pipeline.” However, these queries primarily retrieved studies unrelated to ML pipeline design and yielded no relevant work on PTM reuse pipelines; therefore, they were excluded from the final keyword set. We restricted the search to the 2018–2024 period to capture the emergence and rapid adoption of large-scale PTMs following the introduction of BERT.
The search was conducted across IEEE Xplore, ACM Digital Library, and Google Scholar (queried using the Publish or Perish tool [^5] See https://harzing.com/resources/publish-or-perish). The initial search returned 900 papers. After removing duplicates and non-archival proceedings entries (e.g., editorials, workshop summaries, and short front-matter without full methodological descriptions), 491 papers remained. We excluded such entries because they typically lack sufficient technical detail to analyze PTM reuse pipelines and often overlap with full peer-reviewed publications. The first author then conducted title/abstract screening. Papers were excluded if they focused solely on algorithms without pipeline context, discussed PTMs conceptually without a workflow perspective, or addressed DevOps/governance rather than model-level reuse. Ambiguous cases were discussed with co-authors until consensus. After removing duplicates and irrelevant papers based on titles and abstracts, 300 papers are selected for full-text review.
Our review finds no prior work provides a taxonomy of PTM reuse pipelines or their stages. Most papers describe isolated PTM reuse activities (e.g., fine-tuning, prompt engineering) without situating them in a unified pipeline structure. We developed our taxonomy through open coding. We started with a preliminary API dictionary from prior ML pipeline work and PTM frameworks (Hugging Face, PyTorch), identifying characteristic functions like from_pretrained(), train(), evaluate(). We used contextual cues (mode switches, optimizer presence, output usage) to resolve ambiguities where the same API appears in multiple stages. The first author developed the initial taxonomy, and all authors participated in three rounds of discussion to refine stage definitions. Based on our analysis, we manually identify a minimal, reusable set of PTM-centric stages that consistently appear across both the literature and open-source projects. We focus specifically on model-level, code-observable practices, and intentionally exclude full-stack, DevOps, or governance frameworks (Fitsilis et al., 2024). Table 3 lists the eight identified stages. Stages such as data processing and post-processing are treated at a coarse level, as their fine-grained substeps (e.g., data acquisition, data preparation) vary widely and are often implicit or domain-specific.
@@
The labeling process consisted of two steps. In the first step, the first author labeled all projects, and all authors participated in three rounds of discussion to resolve ambiguous cases and refine the stage definitions. In the second step, we conducted a cross-check by having one co-author independently annotate a randomly selected subset of 282 PTMs based on the finalized stage definitions, without access to the first author’s labels. This subset is generated using a random sampling approach calculated to achieve a 95% confidence level and a 5% margin of error, ensuring it is statistically representative of the full dataset (i.e., 1,059 pipeline instances), resulting in an inter-rater agreement score of 0.81, indicating substantial agreement (Viera et al., 2005).
@@ 4.2.2 Results
Table 3: Unified taxonomy of PTM reuse pipeline stages from our empirical study of 401 projects, arranged chronologically from model initialization to final delivery. Stages that differ from or expand on prior literature are highlighted in bold, including Adaptation (ADPT), Prompting (PROM), Post-Processing (POST), and the newly surfaced Evaluation (EVAL) and Delivery (DLV) stages. Literature-based highlights and our taxonomy-based descriptions are provided for each stage.
@@
Prompting (PROM) Crafting prompts—manually or automatically—to guide generative PTMs (Ramachandran, 2023). Creates prompts manually, templated, or automatically using prompt engineering techniques.
@@
  1. 1.
    Evaluation (EVAL) assesses model outputs systematically using quantitative metrics (e.g., FID score, BLEU, accuracy) or qualitative criteria to measure quality and task alignment. Note that we introduce Evaluation (EVAL) as a distinct stage in the PTM reuse pipeline (Table. 3), even though conventional ML pipelines (Fig. 3) also include an evaluation step, because the two evaluation stages differ in purpose and operational context. In conventional ML pipelines, evaluation typically refers to validation or testing during model development, where models trained from scratch are assessed to guide improvement or selection prior to deployment. In contrast, the Evaluation (EVAL) in PTM reuse pipeline denotes post-inference assessment of a ready-to-use model (an existing PTM or adapted/fine-tuned PTM)’s outputs (e.g., quality, safety, or consistency) during application use rather than during model development.
  2. 2.
@@
Using PTM weights Pretrained weights are used to initialize a custom model structure. This involves mapping parameters onto a new architecture.
Example: Transferring ResNet-50 weights to a custom backbone for human mesh recovery.
Addition New structural components are appended to the PTM backbone to extend its functionality.
@@
In addition, several existing stages take on different emphases in practice. For example, Adaptation appears prominently in real-world pipelines, encompassing not only the addition of a head but also head replacement, selective layer use, architectural changes, and configuration tweaks (Table 4). Fine-tuning itself spans a spectrum of strategies, from fully updating all parameters to partially freezing layers (e.g., freezing lower-level feature extractors while training the higher-level layers), thereby balancing computational efficiency with task specialization. Likewise, Prompting and Post-Processing often serve as critical operational stages in generative tasks, with the latter ensuring output quality and safety.
@@
Feature ExtractionGenerativeDiscriminativeModel InitializationModel InitializationModel InitializationModel AdaptationModel AdaptationModel AdaptationData ProcessingData ProcessingData ProcessingPromptFeature EngineeringFeature EngineeringFeature EngineeringFine-TuningFine-TuningFine-TuningInferenceInferenceInferenceORORORPost-ProcessingPost-ProcessingPost-ProcessingEvaluationEvaluationDeliveryDeliveryLegend:MandatoryOptionalOr ($\geq$1)
Figure 9: Feature model of the PTM Reuse Pipeline variability. Mandatory stages (solid) are shared across all pipeline types. Dashed boxes indicate optional stages. Grey boxes denote an or-group, i.e., at least one of Fine-Tuning or Inference must be present. Cross-tree constraints capture stage availability per pipeline type.
@@
    Generative-Oriented Pipeline. This organization centers on generating novel outputs such as text, code, or images. After model initialization, adaptation, and data processing, a defining stage is prompt, which guides the model’s generative behavior. Depending on the application, the PTM may be used as-is via inference, or fine-tuned for greater task alignment. Feature engineering may be incorporated in hybrid methods (e.g., retrieval-augmented generation). As in other organizations, optional stages such as post-processing and delivery serve to clean, format, or serve the outputs.
@@
Across our dataset, each PTM is associated with one of these pipeline organizations. Feature extraction–oriented pipelines are the most common, accounting for 59% of PTM pipelines, followed by discriminative-oriented pipelines at 21%, and generative-oriented pipelines at 20%.
@@
RQ2: What are the stages and their organization in the reuse pipeline of PTM-based open source projects? Our analysis reveals that PTM reuse pipelines follow three dominant organizations: feature extraction, generative, and discriminative. Across all types, inference is the most consistently used stage, underscoring its central role in practical deployments. While some applications adopt minimal adaptation, complex workflows incorporate full-stage pipelines, including data processing, fine-tuning, post-processing, and delivery. The Model Adaptation stage is especially diverse; developers reuse PTMs as-is, initialize with pre-trained weights, add task-specific heads, or selectively modify layers. These patterns demonstrate that PTM reuse is rarely plug-and-play; instead, customization is key to optimizing performance and aligning with task-specific requirements across varied real-world scenarios.